OpenStudy (anonymous):

What is the value of expression? 5^−8 • 5^−4 A. 5^32 B. 54 C. 5^−4 D. 5^−12

Nnesha (nnesha):

when we multiply same bases we should add their exponents so just add exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

OpenStudy (anonymous):

ok

Nnesha (nnesha):

so what did you get ? \[\large\rm 5^{-8} \times 5^{-4} =5^{??}\] ?

OpenStudy (anonymous):

5^14

OpenStudy (anonymous):

woops i mean 5^12

OpenStudy (anonymous):

sorry

Nnesha (nnesha):

both exponents are negative so -8-4 = ?